OpenStudy (anonymous):

What is the value of p in the equation 4p + 3(p + 4) = 12p + 12? A. 12 B. 1 C. 0 D. -1

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Preetha @Somy

OpenStudy (anonymous):

4p+3p+12=12p 7p+12=12p Subtract 7p from each side 12=5p now just divide by 5

OpenStudy (anonymous):

so its one ? @silenceforthedead

OpenStudy (anonymous):

4p+3p+12=12p+12 7p+12=12p + 12 Subtract 7p and 12 from each side 0=5p Now what times 5 gives you 0

OpenStudy (anonymous):

Sorry i forgot the +12 on my first solution. Look at my second comment

OpenStudy (anonymous):

the answer is c ?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Did you understand how i got to the answer?